4x+5y=6
6x-7y=-20

Answer by stanbon(75887)

4x+5y=6
6x-7y=-20
----------
Multiply thru 1st equation by 6
Multiply thru 2nd equation by 4
----------
24x + 30y = 36
24x - 28y = -80
-----
Subtract and solve for "y":
58y = 116
y = 2
---
solve for "x":
4x+5y = 6
4x + 5*2 = 6
4x = -4
x = -1
-------------------
Cheers,
Stan H.
===================
Cheers,
Stan H.
===================